OpenStudy (anonymous):

simplify this expression: 12x^2y^3 over 3xy

OpenStudy (anonymous):

you have to add the values to make it eaiser

OpenStudy (whpalmer4):

That's\[\frac{12x^2y^3}{3xy}\]

OpenStudy (jdoe0001):

\(\bf \cfrac{12x^2y^3 }{3xy}\implies \cfrac{4\cdot 3\cdot x\cdot x\cdot y\cdot y\cdot y}{3\cdot x\cdot y}\)

OpenStudy (jdoe0001):

anything we can cancel out from there? \(\bf \cfrac{12x^2y^3 }{3xy}\implies \cfrac{4\cdot {\color{red}{ 3\cdot x}}\cdot x\cdot {\color{red}{ y}}\cdot y\cdot y}{{\color{red}{ 3\cdot x\cdot y}}}\quad ?\)

OpenStudy (anonymous):

x,y,3

OpenStudy (whpalmer4):

\[\frac{12x^2y^3}{3xy} = \frac{4 * \cancel{3} * \cancel{x} * x * \cancel{y} * y * y}{\cancel{3} * \cancel{x} * \cancel{y}} = \]